\begin{thebibliography}{10}

\bibitem{amos2017input}
Brandon Amos, Lei Xu, and J~Zico Kolter.
\newblock Input convex neural networks.
\newblock In {\em International conference on machine learning}, pages
  146--155. PMLR, 2017.

\bibitem{arora2018understanding}
Raman Arora, Amitabh Basu, Poorya Mianjy, and Anirbit Mukherjee.
\newblock Understanding deep neural networks with rectified linear units.
\newblock In {\em International Conference on Learning Representations}, 2018.

\bibitem{averkov2025on}
Gennadiy Averkov, Christopher Hojny, and Maximilian Merkert.
\newblock On the expressiveness of rational relu neural networks with bounded
  depth.
\newblock In {\em The Thirteenth International Conference on Learning
  Representations}, 2025.

\bibitem{bakaev2025depth}
Egor Bakaev, Florestan Brunck, Christoph Hertrich, Daniel Reichman, and Amir
  Yehudayoff.
\newblock On the depth of monotone relu neural networks and icnns.
\newblock {\em arXiv preprint arXiv:2505.06169}, 2025.

\bibitem{bakaev2025better}
Egor Bakaev, Florestan Brunck, Christoph Hertrich, Jack Stade, and Amir
  Yehudayoff.
\newblock Better neural network expressivity: subdividing the simplex.
\newblock {\em arXiv preprint arXiv:2505.14338}, 2025.

\bibitem{bunning2021input}
Felix B{\"u}nning, Adrian Schalbetter, Ahmed Aboudonia, Mathias~Hudoba
  de~Badyn, Philipp Heer, and John Lygeros.
\newblock Input convex neural networks for building mpc.
\newblock In {\em Learning for dynamics and control}, pages 251--262. PMLR,
  2021.

\bibitem{chen2018optimal}
Yize Chen, Yuanyuan Shi, and Baosen Zhang.
\newblock Optimal control via neural networks: A convex approach.
\newblock In {\em International Conference on Learning Representations}, 2019.

\bibitem{cover2006elements}
Thomas Cover and Joy Thomas.
\newblock {\em Elements of information theory}.
\newblock John Wiley \& Sons, 2006.

\bibitem{daubechies2022neural}
Ingrid Daubechies, Ronald DeVore, Nadav Dym, Shira Faigenbaum-Golovin, Shahar~Z
  Kovalsky, Kung-Chin Lin, Josiah Park, Guergana Petrova, and Barak Sober.
\newblock Neural network approximation of refinable functions.
\newblock {\em IEEE Transactions on Information Theory}, 69(1):482--495, 2022.

\bibitem{daubechies2022nonlinear}
Ingrid Daubechies, Ronald DeVore, Simon Foucart, Boris Hanin, and Guergana
  Petrova.
\newblock Nonlinear approximation and (deep) relu networks.
\newblock {\em Constructive Approximation}, 55(1):127--172, 2022.

\bibitem{devore2021neural}
Ronald DeVore, Boris Hanin, and Guergana Petrova.
\newblock Neural network approximation.
\newblock {\em Acta Numerica}, 30:327--444, 2021.

\bibitem{eldan2016power}
Ronen Eldan and Ohad Shamir.
\newblock The power of depth for feedforward neural networks.
\newblock In {\em 29th Annual Conference on Learning Theory}, volume~49, pages
  907--940. PMLR, 2016.

\bibitem{gagneux2025convexity}
Anne Gagneux, Mathurin Massias, Emmanuel Soubies, and R{\'e}mi Gribonval.
\newblock Convexity in relu neural networks: beyond icnns?
\newblock {\em Journal of Mathematical Imaging and Vision}, 67(4):40, 2025.

\bibitem{grillo2025depth}
Moritz Grillo, Christoph Hertrich, and Georg Loho.
\newblock Depth-bounds for neural networks via the braid arrangement.
\newblock {\em arXiv preprint arXiv:2502.09324}, 2025.

\bibitem{grunbaum2003convex}
Branko Grünbaum.
\newblock {\em Convex Polytopes}.
\newblock Springer, 2003.

\bibitem{haase2023lower}
Christian~Alexander Haase, Christoph Hertrich, and Georg Loho.
\newblock Lower bounds on the depth of integral relu neural networks via
  lattice polytopes.
\newblock In {\em International Conference on Learning Representations}, 2023.

\bibitem{hertrich2023towards}
Christoph Hertrich, Amitabh Basu, Marco Di~Summa, and Martin Skutella.
\newblock Towards lower bounds on the depth of relu neural networks.
\newblock {\em SIAM Journal on Discrete Mathematics}, 37(2):997--1029, 2023.

\bibitem{lemaire2024new}
Vincent Lemaire, Christian Yeo, et~al.
\newblock A new input convex neural network with application to options
  pricing.
\newblock {\em arXiv preprint arXiv:2411.12854}, 2024.

\bibitem{liang2017why}
Shiyu Liang and R.~Srikant.
\newblock Why deep neural networks for function approximation?
\newblock In {\em International Conference on Learning Representations}, 2017.

\bibitem{maclagan2021introduction}
Diane Maclagan and Bernd Sturmfels.
\newblock {\em Introduction to tropical geometry}, volume 161.
\newblock American Mathematical Society, 2021.

\bibitem{mukherjee2017lower}
Anirbit Mukherjee and Amitabh Basu.
\newblock Lower bounds over boolean inputs for deep neural networks with relu
  gates.
\newblock {\em arXiv preprint arXiv:1711.03073}, 2017.

\bibitem{telgarsky2016benefits}
Matus Telgarsky.
\newblock Benefits of depth in neural networks.
\newblock In {\em 29th Annual Conference on Learning Theory}, volume~49, pages
  1517--1539. PMLR, 2016.

\bibitem{west2001introduction}
Douglas~Brent West et~al.
\newblock {\em Introduction to graph theory}, volume~2.
\newblock Prentice hall Upper Saddle River, 2001.

\bibitem{xing2024optimization}
Shanshuo Xing, Jili Zhang, and Song Mu.
\newblock An optimization-oriented modeling approach using input convex neural
  networks and its application on optimal chiller loading.
\newblock In {\em Building Simulation}, volume~17, pages 639--655. Springer,
  2024.

\bibitem{yarotsky2017error}
Dmitry Yarotsky.
\newblock Error bounds for approximations with deep relu networks.
\newblock {\em Neural Networks}, 94:103--114, 2017.

\bibitem{zaslavsky1975facing}
Thomas Zaslavsky.
\newblock {\em Facing up to arrangements: Face-count formulas for partitions of
  space by hyperplanes: Face-count formulas for partitions of space by
  hyperplanes}, volume 154.
\newblock American Mathematical Soc., 1975.

\bibitem{zhang2018tropical}
Liwen Zhang, Gregory Naitzat, and Lek-Heng Lim.
\newblock Tropical geometry of deep neural networks.
\newblock In {\em International Conference on Machine Learning}, pages
  5824--5832. PMLR, 2018.

\bibitem{ziegler1995lectures}
Günter~M. Ziegler.
\newblock {\em Lectures on Polytopes}.
\newblock Springer, 1995.

\end{thebibliography}

\documentclass{article}  

\usepackage{amssymb, amsmath, amsthm}
\usepackage[hidelinks]{hyperref}
\usepackage[capitalize, noabbrev]{cleveref}
\usepackage[nocompress]{cite}

\usepackage{graphicx}

\newtheorem{lemma}{Lemma}
\newtheorem{proposition}{Proposition}
\newtheorem{theorem}{Theorem}

\newcommand{\R}{\mathbb{R}}
\newcommand{\N}{\mathbb{N}}

\usepackage{enumitem}
\crefalias{enumi}{theorem}
\crefalias{enumii}{theorem}
\crefalias{enumiii}{theorem}

\newcommand{\maxb}[1]{\max\{#1\}}
\newcommand{\conv}[1]{\mathrm{conv}(#1)}
\newcommand{\nn}[1]{\mathsf{ReLU}_n(#1)}
\newcommand{\cpwl}{\mathsf{CPWL}_n}
\newcommand{\lm}{\mathsf{LM}_n}

\title{On Minimal Depth in Neural Networks}

\author{Juan L. valerdi}
\date{}

\begin{document} 
\maketitle

\begin{abstract}
  Understanding the relationship between the depth of a neural network and its representational capacity is a central problem in deep learning theory. In this work, we develop a geometric framework to analyze the expressivity of ReLU networks with the notion of depth complexity for convex polytopes. The depth of a polytope recursively quantifies the number of alternating convex hull and Minkowski sum operations required to construct it. This geometric perspective serves as a rigorous tool for deriving depth lower bounds and understanding the structural limits of deep neural architectures.

  We establish lower and upper bounds on the depth of polytopes, as well as tight bounds for classical families. These results yield two main consequences. First, we provide a purely geometric proof of the expressivity bound by Arora et al. (2018), confirming that $\lceil \log_2(n+1)\rceil$ hidden layers suffice to represent any continuous piecewise linear (CPWL) function. Second, we prove that, unlike general ReLU networks, convex polytopes do not admit a universal depth bound. Specifically, the depth of cyclic polytopes in dimensions $n \geq 4$ grows unboundedly with the number of vertices. This result implies that Input Convex Neural Networks (ICNNs) cannot represent all convex CPWL functions with a fixed depth, revealing a sharp separation in expressivity between ICNNs and standard ReLU networks.
\end{abstract}

\section{Introduction}\label{sec:intro}

A \textit{ReLU neural network} of depth $m$, or with $m$ hidden layers, is a function $f:\R^n\rightarrow\R$ expressed as
\begin{equation*} 
    f = T^{(m+1)} \circ \sigma_m \circ T^{(m)} \circ \sigma_{m-1} \circ \dots \circ T^{(2)} \circ \sigma_1 \circ T^{(1)},
\end{equation*}
where $T^{(l)}:\R^{n_{l-1}} \rightarrow \R^{n_l}$ are $m + 1$ affine transformations, $n_0 = n$ and $n_{m+1} = 1$. The activation function $\sigma_l$ corresponds to the vectorized ReLU function $\sigma_l(x_1, \ldots, x_{n_l}) = (\maxb{x_1, 0}, \maxb{x_2, 0}, \ldots, \maxb{x_{n_l}, 0})$.

ReLU neural networks are continuous piecewise linear (CPWL) functions \cite{arora2018understanding, devore2021neural}. An important problem in learning theory is to characterize which CPWL functions can be represented by a ReLU network of a given depth. Such expressivity results are relevant for quantifying depth-dependent convergence rates for various function classes \cite{daubechies2022neural, daubechies2022nonlinear, devore2021neural, yarotsky2017error} and enhancing our theoretical understanding of deep architectures \cite{arora2018understanding, telgarsky2016benefits, eldan2016power, liang2017why}.

We denote the set of all CPWL functions from $\R^n\rightarrow\R$ as $\cpwl$, and the subset of these functions that can be represented by ReLU networks of depth $m$ as $\nn{m}$.

It has been shown that $\lceil \log_2(n + 1) \rceil$ hidden layers are sufficient to represent any CPWL function.
\begin{theorem}[Arora, Basu, Mianjy and Mukherjee \cite{arora2018understanding}, Theorem 2.1] \label{thm:arora}
  \[
    \cpwl = \nn{\lceil \log_2(n + 1) \rceil}.
  \]
\end{theorem}

The minimal depth required to represent any CWPL function remains an open question. Hertrich, Basu, Di Summa, and Skutella \cite{hertrich2023towards} proposed an approach that reduces the problem to analyzing a single function.

\begin{proposition}[Hertrich, Basu, Di Summa and Skutella \cite{hertrich2023towards}, Proposition 1.6] \label{prop:maxn}
  \[
    \cpwl = \nn{m} \Longleftrightarrow \maxb{x_1, x_2, \ldots, x_n, 0} \in \nn{m}.
  \]
\end{proposition}

Following \cref{prop:maxn}, the goal is to find the minimum $m$ such that the function $\maxb{x_1, \ldots, x_n, 0}\in\nn{m}$. 

Hertrich, Basu, Di Summa, and Skutella \cite{hertrich2023towards} also conjectured that this depth is exactly $\lceil \log_2(n + 1) \rceil$, which had been proven for dimensions $n \leq 3$ \cite{devore2021neural, mukherjee2017lower}. Much of the progress toward resolving this problem has come from formulating \cref{prop:maxn} in the language of convex polytopes, i.e. the convex hull of finitely many points \cite{grunbaum2003convex, ziegler1995lectures}. Haase, Hertrich, and Loho \cite{haase2023lower} proved the conjecture for ReLU networks with integer weights. This result was later extended by Averkov, Hojny, and Merkert \cite{averkov2025on} to networks with decimal fractional weights. Furthermore, Grillo, Hertrich, and Loho \cite{grillo2025depth} proved the conjecture for $n = 4$ in networks compatible with specific polyhedral complexes. More recently, the conjecture was disproved by Bakaev, Brunck, Hertrich, Stade, and Yehudayoff \cite{bakaev2025better} via polyhedral subdivisions of the simplex; they showed that $\lceil \log_3(n - 1) \rceil + 1$ hidden layers are sufficient to represent any CPWL function.

In this work, we introduce the \textit{depth complexity} of polytopes, a geometric analogue of neural network depth, to analyze the expressivity of ReLU networks. Although this concept has been used implicitly in prior work \cite{hertrich2023towards,haase2023lower,bakaev2025better}, we provide a systematic study of it.

We denote the depth complexity of a polytope $P$ as $d(P)$ and define it recursively as follows. If $P$ consists of a single point, we define $d(P) = 0$. If $P$ is not a single point, we define $d(P) = m$, where $m$ is the smallest positive integer such that
\begin{equation} \label{eq:depth}
  P = \sum_{i = 1}^q \conv{P_{i1}, P_{i2}}, \quad \text{with } d(P_{ij}) < m \quad \forall i, j.
\end{equation}
Here, $\conv{P_1, P_2} = \conv{P_1 \cup P_2}$ and $P_1 + P_2 = \{a_1 + a_2: a_i\in P_i\}$ are the \textit{convex hull} and \textit{Minkowski sum} operations respectively. 

At its core, the depth complexity of a polytope quantifies the number of alternating steps of convex hulls and Minkowski sums required to construct it. This serves as a measure of how ``complex'' it is to represent a polytope using these two operations. Under this measure, the simplest polytope is a point and depth-1 polytopes correspond to zonotopes, which are Minkowski sums of segments.

Before connecting the depth of polytopes and ReLU networks, we introduce some terminology. A \textit{linear max function} is a function of the form $f(x) = \maxb{a_1\cdot x, \ldots, a_p\cdot x}$ with $a_i\in\R^n$ for all $i$. Notably, linear max functions are convex and \textit{positively homogeneous}, meaning they satisfy $f(\lambda x) = \lambda f(x)$ for $\lambda \geq 0$.

Let $\lm$ denote the collection of linear max functions from $\R^n\rightarrow\R$, and let $\mathcal{P}_n$ represent the set of convex polytopes in $\R^n$. These two sets constitute semirings with the \textit{max} and \textit{sum} operations, and \textit{convex hull} and \textit{Minkowski sum} operations, respectively.

For the function $f(x) = \maxb{a_1\cdot x, \ldots, a_p\cdot x}$, its \textit{Newton polytope} is defined as
\[
  \mathcal{N}f := \conv{a_1, \ldots, a_p}.
\]
Similarly, for the polytope $P = \conv{a_1, \ldots, a_p}$, the associated \textit{support function} is
\[
  \mathcal{F}P(x) := \maxb{a_1\cdot x, \ldots, a_p\cdot x}.
\]
The mappings $\mathcal{N}$ and $\mathcal{F}$ are isomorphisms between the semirings ($\lm$, max, $+$) and ($\mathcal{P}_n$, conv, $+$) \cite{maclagan2021introduction, zhang2018tropical}. 

The following result establishes a depth connection between ReLU networks and polytopes.

\begin{theorem}[Hertrich, Basu, Di Summa and Skutella \cite{hertrich2023towards}, Theorem 5.2] \label{thm:relu-poly}
  Let $f\in\cpwl$ be a positively homogeneous function. Then, $f\in\nn{m}$ if and only if there exist $f_i\in \lm, i = 1,2, $ such that $f = f_1 - f_2$ and $d(\mathcal{N}f_i) \leq m$.
\end{theorem}

Combining \cref{prop:maxn} with \cref{thm:relu-poly} gives an equivalent geometric statement of \cref{prop:maxn}.

\begin{theorem} \label{thm:cpwl-poly}
  \[
    \cpwl = \nn{m}
  \]
  if and only if there exist polytopes $P_i, i = 1,2, $ such that
  \[
    P_1 = \mathcal{N}(\maxb{x_1, x_2, \ldots, x_n, 0}) + P_2 \quad \text{and} \quad d(P_i) \leq m.
  \]
\end{theorem}

To fully leverage \cref{thm:cpwl-poly}, it is relevant to establish foundational results on the depth complexity of polytopes. In addition, two natural questions arise: What is the minimum depth required to represent any polytope? And, what is the depth complexity of simplices like $\mathcal{N}(\maxb{x_1, x_2, \ldots, x_n, 0})$?\\

\textit{Depth Complexity Results}\\

In \cref{sec:upper}, we derive depth complexity upper bounds based on Minkowski sums, affine transformations, convex hull representations, and combinatorial data such as the number of vertices, edges, and 2-faces. Combined with \cref{thm:relu-poly}, these bounds yield depth estimates for ReLU network representations of linear max functions.

An affine transformation is a mapping $\varphi(x) = Ax + b$, where $A$ is a linear map and $b$ is a translation vector. A convex hull representation of a polytope $P$ in terms of other polytopes $P_1, \ldots, P_k$ is a decomposition of the form $P = \conv{P_1, P_2, \ldots, P_k}$. 

If the polytopes $P_i$ are the vertices, edges or 2-faces of $P$, we can compute depth estimates based on their counts. For a polytope with $k$ vertices, we obtain the upper bound $\lceil \log_2 k\rceil$, which is tight for certain families of polytopes.

In \cref{sec:lower}, we derive lower bounds on depth complexity using its graph and face structure. Let $G(P)$ denote the 1-skeleton of $P$, i.e. the graph whose vertices are those of $P$ and in which two vertices are adjacent when they span an edge of $P$.

We show that if $G(P)$ contains a complete subgraph on $k$ vertices, then $d(P) \geq \lceil \log_2 k\rceil$. This follows from the fact that complete subgraphs are pervasive in expressions like (\ref{eq:depth}): if $G(P)$ contains a complete subgraph, then at least a summand in the Minkowski sum will have one, which then propagates through the polytopes in the convex hull. 

Moreover, for any face $F$ of $P$, we have $d(F) \leq d(P)$. Hence, $d(P)$ is at least the maximum depth complexity among its faces. 

In \cref{sec:polytopes}, we determine the depth complexity of several families of polytopes, including pyramids, bipyramids, prisms, simplices, and cyclic polytopes. Using depth estimates based on vertex counts and complete subgraphs, it follows directly that polytopes with $k$ vertices whose graphs are complete, such as simplices and cyclic polytopes, have depth complexity $\lceil \log_2 k\rceil$.

This result answers both motivating questions posed earlier. For simplices, the main application follows from \cref{thm:cpwl-poly}, where we recover the same bound as in \cref{thm:arora}, thus providing an alternative geometric proof. In contrast, for dimension $n\geq 4$, cyclic polytopes form a family whose depth grows unbounded as the number of vertices increases. This shows that there is no universal upper depth bound to represent polytopes--unlike the case for ReLU networks in \cref{thm:arora}. Notably, the proofs of these results rely only on elementary polyhedral theory.

We conclude \cref{sec:polytopes} by showing that cyclic polytopes not only exhibit unbounded depth, but in dimensions $n\geq 5$, their Minkowski sum with zonotopes yield, for any given depth $m$, families of polytopes with an increasing number of vertices, all of depth $m$.\\

\textit{Implications for Input Convex Neural Networks}\\

Input Convex Neural Networks (ICNNs) \cite{amos2017input} are ReLU networks constrained to use only monotone affine transformations and skip connections from the input layer. Here, a monotone affine transformation means an affine map $Ax + b$ where the matrix $A$ has nonnegative entries.

ICNNs are interesting because they can represent any convex CPWL function \cite{bakaev2025depth, gagneux2025convexity} and have found many  successful applications \cite{lemaire2024new, xing2024optimization, bunning2021input, chen2018optimal}.

Following \cite{bakaev2025depth}, we can define under the ICNN model, a corresponding depth complexity $d_0(P)$ for a polytope $P$: $d_0(P) = 0$ if $P$ is a single point, else $d_0(P) = m$ for $m$ the minimum value such that
\[
  P = \sum_{i = 1}^p \conv{P_{i1}, P_{i2}}, \quad \text{with } d_0(P_{i1}) < m \text{ and } d_0(P_{i2}) = 0 \quad \forall i.
\]

In \cite{bakaev2025depth} the authors showed that $d_0(\mathcal{N}(\maxb{x_1, \ldots, x_n})) = n$, and that for $n=3$ there exists a family of polytopes with unbounded $d_0$-complexity. For ICNNs, $d_0(P)$ also constitutes the minimum ICNN depth required to represent $\mathcal{F} P$.

Since $d(P) \leq d_0(P)$, and cyclic polytopes with $k$ vertices satisfy $d(P) = \lceil \log_2 k \rceil$ for $n \geq 4$, they also form an unbounded depth family for $d_0$. Thus, although ICNNs can represent every convex CPWL function, no fixed depth bound suffices in general. This contrasts with the universal depth bound for ReLU networks established in \cref{thm:arora}.\\

\textbf{Acknowledgements.} I am deeply grateful to Francisco Santos for \cref{thm:graph}, valuable discussions, and his hospitality at the University of Cantabria. I also thank Ansgar Freyer for discussions on the depth complexity of simplices.

\section{Depth Upper Bounds} \label{sec:upper}

We begin by computing basic bounds for Minkowski sums, convex hulls and affine maps. 

\begin{proposition}\label{prop:basic_upper} 
  Let $P_i, i = 1, 2,$ be polytopes with $d(P_i) \leq m_i$ and $\varphi$ be an affine transformation. Then,
  \begin{enumerate}[label=(\alph*)]
  \item $d(P_1 + P_2) \leq \maxb{m_1, m_2}$
  \item $d(\conv{P_1, P_2}) \leq \maxb{m_1, m_2} + 1$
  \item $d(\varphi(P_1)) \leq d(P_1)$
  \end{enumerate}
\end{proposition}
\begin{proof}
  These depth bounds follow immediately from the depth complexity definition. If $d(P_i) \leq m_i$, it implies that we can express $P_i$ as 
  \[
    P_i = \sum_{j = 1}^{q_i} \conv{Q_{j, i}, R_{j, i}},\quad d(Q_{j, i}), d(R_{j, i}) < m_i \quad \forall i, j.
  \]
  Then,
  \[
    P_1 + P_2 = \sum_{j = 1}^{q_1} \conv{Q_{j, 1}, R_{j, 1}} + \sum_{j = 1}^{q_2} \conv{Q_{j, 2}, R_{j, 2}}.
  \]
  All $Q_{j, i}$ and $P_{j, i}$ have depth strictly less than $\maxb{m_1, m_2}$, therefore $d(P_1 + P_2) \leq \maxb{m_1, m_2}$.

  Similarly, since $d(P_i) \leq \maxb{m_1, m_2}$, it follows by definition that $$d(\conv{P_1, P_2}) \leq \maxb{m_1, m_2} + 1.$$
  Let $\varphi(x) = Ax + b$ be an affine map, if $d(P_1) = 0$, it is trivial that $d(\varphi(P_1)) = 0$. For the purpose of induction, assume the statement is true up to $m - 1$. If $d(P_1) = m$, then
  \[
    \varphi(P_1) = \varphi\Big(\sum_{j = 1}^{q_1} \conv{Q_{j, 1}, R_{j, 1}}\Big) = \sum_{j = 1}^{q_1} \conv{AQ_{j, 1}, AR_{j, 1}} + b, 
  \]
  and by induction hypothesis $d(AQ_{j, 1}), d(AR_{j, 1}) < m$, implying $d(\varphi(P_1)) \leq d(P_1)$.
\end{proof}

If an affine transformation $\varphi$ is invertible, note that $d(\varphi(P)) = d(P)$. This also implies that $d(P_1 + P_2) = d(P_1)$ for $d(P_2) = 0$, which is simply a translation.

When $P = \conv{P_1, \ldots, P_k}$, iterating \cref{prop:basic_upper}(b) yields an upper bound for $d(P)$. There are several ways to iterate a binary operation on $k$ elements, corresponding to different choices of binary bracketing.

Any binary bracketing of $\conv{P_1, \ldots, P_k}$ can be encoded by a full binary tree whose leaves are labeled by $1, \ldots, k$: each internal node records one application of $\conv{\cdot,\cdot}$, and its two children correspond to the two subexpressions being combined. Conversely, every full binary tree with leaves labeled by $1, \ldots, k$ determines a binary bracketing of $\conv{P_1, \ldots, P_k}$. For a leaf $i$, let $\ell_i$ denote its depth in the tree, that is, its distance to the root.

We identify the smallest bound obtainable by iterating \cref{prop:basic_upper}(b) using Kraft's inequality \cite[Chapter 5.2]{cover2006elements}. 

\begin{theorem}[Kraft inequality for binary trees]
Let $\ell_1, \ldots, \ell_k$ be positive integers. Then, there exists a binary tree with leaves at depths $\ell_1, \ldots, \ell_k$ if and only if
\[
  \sum_{i=1}^k 2^{-\ell_i} \le 1.
\]
\end{theorem}

\begin{proposition} \label{prop:optimal_hull}
  Let $P = \conv{P_1, \ldots, P_k}$ with $d(P_i) \leq m_i$ for all $i$. Then
  \[
    d(P) \leq \Big\lceil \log_2\Big(\sum_{i=1}^k 2^{m_i}\Big)\Big\rceil.
  \]
  Moreover, this is the smallest bound that can be obtained by iterating \cref{prop:basic_upper}(b) over the polytopes $P_i$.
\end{proposition}
\begin{proof}
  For $k = 1$, the statement is immediate. Assume $k \geq 2$, fix a binary bracketing of $\conv{P_1, \ldots, P_k}$, and let $T$ be the corresponding full binary tree.

  We claim that the upper bound produced by iterating \cref{prop:basic_upper}(b) according to $T$ is
  \[
    M_T = \max_{1 \leq i \leq k}(m_i + \ell_i).
  \]
  We prove this by induction on the number of leaves of $T$. If $T$ has two leaves, the result is a direct consequence of \cref{prop:basic_upper}(b). If $T$ has more than two leaves, it means that the root has two subtrees $T_1$ and $T_2$, and applying \cref{prop:basic_upper}(b) at the root gives
  \[
    M_T = \maxb{M_{T_1}, M_{T_2}} + 1.
  \]
  By induction, each $M_{T_j}$ is the maximum of $m_i + \ell_i - 1$ over the leaves of $T_j$, because their depths inside the subtree are one less than in $T$. Hence
  \[
    M_T = \max_{1 \leq i \leq k}(m_i + \ell_i).
  \]

  Every binary tree satisfies Kraft's inequality
  \[
    \sum_{i=1}^k 2^{-\ell_i} \leq 1.
  \]
  Therefore,
  \[
    m_i + \ell_i \leq M_T \quad \forall i
  \]
  implies
  \[
    2^{m_i - M_T} \leq 2^{-\ell_i} \quad \forall i,
  \]
  and summing over $i$, we obtain
  \[
    2^{-M_T}\sum_{i=1}^k 2^{m_i} \leq 1,
  \]
  so
  \begin{equation}
    \label{eq:kraft}
    M_T \geq \Big\lceil \log_2\Big(\sum_{i=1}^k 2^{m_i}\Big)\Big\rceil.
  \end{equation}

  To show that this lower bound is attained for some $T$, set
  \[
    M = \Big\lceil \log_2\Big(\sum_{i=1}^k 2^{m_i}\Big)\Big\rceil
    \qquad\text{and}\qquad
    L_i = M - m_i.
  \]
  Therefore,
  \[
    \sum_{i=1}^k 2^{-L_i} = 2^{-M}\sum_{i=1}^k 2^{m_i} \leq 1.
  \]
  By Kraft's inequality, there exists a binary tree $T_0$ with leaves at depths $L_1, \ldots, L_k$. If $T_0$ is not full, contract every internal node with exactly one child. This produces a full binary tree $T$ with the same labeled leaves, and the depth $\ell_i$ of leaf $i$ in $T$ satisfies $\ell_i \leq L_i$ for all $i$. The tree $T$ therefore determines a binary bracketing of $\conv{P_1, \ldots, P_k}$, and for this bracketing
  \[
    M_T = \max_{1 \leq i \leq k}(m_i + \ell_i) \leq \max_{1 \leq i \leq k}(m_i + L_i) = M.
  \]
  Combined with (\ref{eq:kraft}), we get $M_T = M$, which is then the smallest upper bound obtainable by iterating \cref{prop:basic_upper}(b).
\end{proof}

If $m_i \leq m$ for all $i$ in \cref{prop:optimal_hull}, then
\[
  d(P) \leq \Big\lceil \log_2\Big(\sum_{i=1}^k 2^{m_i}\Big)\Big\rceil \leq \lceil \log_2 k \rceil + m.
\]
We will use this simplified form to obtain depth estimates when the numbers of low-dimensional faces of a polytope are known.

For this, we also need the following concepts: $\rho(P)$ and $\rho_2(P)$ are the minimum number of edges and 2-faces, respectively, needed to cover all vertices of the polytope $P$; $\nu(P)$ is the size of a maximum matching of the graph $G(P)$, i.e. the largest set of edges in $G(P)$ with no common vertices; $P_x$ is the vertex figure of $P$ at a vertex $x$, i.e. the intersection of $P$ with a hyperplane that separates $x$ from the other vertices of $P$.

\begin{proposition} \label{prop:f}
  Let $f_0, f_1, f_2$ be the number of vertices, edges and 2-faces respectively of a polytope $P$. Then, we have the following depth bounds depending of each $f_i$ individually:
  \begin{enumerate}[label=(\alph*)]
  \item $\text{For } f_0:\quad d(P) \leq \lceil \log_2 f_0 \rceil$
  \item $\text{For } f_1:\quad d(P) \leq \Big\lceil\log_2\Big(\Big\lfloor\frac{2f_1}{n} - \min\Big(n, \Big\lfloor \frac{1+\sqrt{1+8f_1}}{4} \Big\rfloor\Big)\Big\rfloor\Big)\Big\rceil + 1$
  \item $\text{For } f_2:\quad d(P) \leq \Big\lceil \log_2 \Big(f_2 - \binom{n}{2} + 1\Big)\Big\rceil + 2$
  \end{enumerate}
\end{proposition}
\begin{proof}
  Any polytope $P$ can be expressed as $P = \conv{P_1, \ldots, P_{f_j}}$, where $P_i$ are $j$-dimensional faces of $P$.\\

  (a) If $P_i$ are the vertices, then $d(P_i) = 0$, and by \cref{prop:optimal_hull} we obtain  
  \[
    d(P) \leq \lceil \log_2 f_0 \rceil.
  \]

  (b) If $P_i$ are the edges of $P$, then $d(P_i) = 1$. We do not need all $f_1$ edges as $P$ can be written as  
  \[
    P = \conv{P_{i_1}, \ldots, P_{i_{\rho(P)}}},
  \]  
  where $\rho(P)$ is the size of a minimum edge cover $P_{i_1}, \ldots, P_{i_{\rho(P)}}$. 

  To estimate $\rho(P)$, we use standard relations between $\rho(P)$, $\nu(P)$, $f_0$, and $f_1$ (see \cite[Chapter 3.1]{west2001introduction}):  
  \[
    \rho(P) + \nu(P) = f_0, 
    \qquad 
    \nu(P) \geq \min\big(n, \lfloor f_0/2 \rfloor\big).
  \]  
  Furthermore, the degree sum formula and the complete graph vertices bound yield  
  \[
    nf_0 \leq 2f_1 \leq f_0(f_0 - 1).
  \]  
  Combining these inequalities gives  
  \[
    \rho(P) \leq \Big\lfloor \frac{2f_1}{n} - \min\Big(n, \Big\lfloor \tfrac{1+\sqrt{1+8f_1}}{4} \Big\rfloor\Big)\Big\rfloor,
  \]  
  from which the stated bound on $d(P)$ in terms of $f_1$ follows.
  
  (c) Consider the case of $2$-faces. In \cref{sec:polytopes} we show that the depth complexity of polygons is at most $2$. Hence, similarly as (b), it suffices to bound $\rho_2(P)$, the minimum number of $2$-faces covering all vertices of $P$, and apply \cref{prop:optimal_hull}.  

  Let $P_x$ be the vertex figure of $P$ at a vertex $x$. Since $\dim P_x = n-1$, the degree sum formula implies that $P_x$ has at least $\binom{n}{2}$ edges. These correspond bijectively to the $2$-faces of $P$ incident to $x$. Thus, even after removing any $\binom{n}{2} - 1$ $2$-faces, every vertex of $P$ is still contained in at least one $2$-face. Hence $\rho_2(P) \leq f_2 - \binom{n}{2} + 1$, which yields  
  \[
    d(P) \leq \Big\lceil \log_2 \big(f_2 - \binom{n}{2} + 1\big)\Big\rceil + 2. \qedhere
  \]
\end{proof}

\section{Depth Lower Bounds} \label{sec:lower}

In this section, we study depth lower bounds for polytopes, obtained from their graphs and from the depth complexity of their faces.

\begin{proposition} \label{prop:faces} 
  Let $P$ be a polytope. Then for any nonempty face $F$ of $P$, we have $d(P) \geq d(F)$.
\end{proposition}
\begin{proof}
  If $d(P) = 0$, there is nothing to prove. If $d(P) = 1$, then $P$ is a zonotope, and any face $F$ is also a zonotope, hence $d(F) \leq 1$.

  Now suppose, for induction, that the statement holds for all polytopes of depth at most $m-1$, and assume $d(P) = m$. By definition, we can write
  \[
    P = \sum_{i=1}^q \conv{P_{i1}, P_{i2}}, \quad d(P_{ij}) < m \quad \forall i, j.
  \]
  
  Let $F$ be a face of $P$. Then $F = \sum_{i=1}^q F_i$, where each $F_i$ is a face of $\conv{P_{i1}, P_{i2}}$. This implies that $F_i$ can be written as $F_i = \conv{F_{i1}, F_{i2}}$, where $F_{ij}$ is a (possibly empty) face of $P_{ij}$. 

  By the induction hypothesis, whenever $F_{ij} \neq \emptyset$, we have $d(F_{ij}) < m$. If both $F_{i1}$ and $F_{i2}$ are nonempty, then by \cref{prop:basic_upper}(b) we have $d(F_i) \leq m$. If exactly one of $F_{i1}, F_{i2}$ is nonempty, then $F_i$ is equal to that nonempty face, and hence $d(F_i) \leq m$. 

Therefore, in all cases each $F_i$ satisfies $d(F_i) \leq m$. Finally, since $F = \sum_{i=1}^q F_i$, \cref{prop:basic_upper}(a) implies $d(F) \leq m = d(P)$.
\end{proof}

To obtain a lower bound from complete subgraphs contained in $G(P)$, we first state the following lemma.

\begin{lemma}\label{lem:subgraph}
  If the graph of a polytope $G(P)$ contains a complete subgraph with $k\geq 3$ vertices, and $P$ can be decomposed as $P = \sum_{i=1}^q P_i$, then at least one of $G(P_i)$ also contains a complete subgraph with $k$ vertices.
\end{lemma}
\begin{proof}
  Any edge $e=[x_1,x_2]$ of $P$ can be written as $e = \sum_{i=1}^q e_i$, where each $e_i$ is either a vertex or an edge of $P_i$. If $e_i$ is a vertex, we regard it as a degenerate edge of length zero, so we can say that each edge of $P$ decomposes as a sum of edges of the $P_i$.

  Now let $u,v,w$ be three vertices in the complete subgraph of $G(P)$ (so $[u,v],[u,w],[v,w]$ are edges of $P$). Write
  \[
    [u, v] = \sum_{i=1}^q [u_i, v_i], \quad [u, w] = \sum_{i=1}^q [u_i, w_i], \quad [v, w] = \sum_{i=1}^q [v_i, w_i],
  \]
  where $u=\sum_i u_i$, $v=\sum_i v_i$, $w=\sum_i w_i$ with $u_i,v_i,w_i$ vertices of $P_i$.  

  The edges $[u_i, v_i], [u_i, w_i], [v_i, w_i]$ are parallel to $[u, v], [u, w], [v, w]$ respectively, since otherwise their Minkowski sum would yield a polytope of dimension at least two, contradicting the fact that the corresponding sum is an edge. This implies that each triangle $[u_i, v_i, w_i]$ is homothetic to the triangle $[u, v, w]$, and therefore all edge length ratios are preserved 
  \[
    \frac{|u_i - v_i|}{|u - v|} = \frac{|u_i - w_i|}{|u - w|} = \frac{|v_i - w_i|}{|v - w|}.
  \]
  Note that some triangles $[u_i,v_i,w_i]$ may be degenerate (if two or all three vertices coincide). In this case, the triangle is still homothetic to $[u,v,w]$ with homothety factor $0$, so the ratio identities above remain valid.

  As the triangle $[u, v, w]$ is not degenerate, at least one index $j$ has these ratios nonzero, so $u_j,v_j,w_j$ are distinct vertices forming a triangle in $G(P_j)$.  

  For any other vertex $z$ in the clique of $G(P)$, applying the same argument to the triangles $[u,v,z]$ and $[u, w, z]$ shows that the same index $j$ yields non-degenerate homothetic triangles $[u_j, v_j, z_j]$ and $[u_j, w_j, z_j]$, with $z_j$ distinct from $u_j, v_j, w_j$ and adjacent to them. Repeating this for every vertex in the clique, we conclude that $G(P_j)$ contains a complete subgraph with $k$ vertices.
\end{proof}

\begin{theorem}\label{thm:graph}
  If the graph of a polytope $G(P)$ contains a complete subgraph with $k\geq 3$ vertices, then $d(P) \geq \lceil\log_2 k\rceil$.
\end{theorem}
\begin{proof}
  Suppose a subgraph of $G(P)$ is complete and contains $k = 3$ or $k = 4$ vertices. If we assume $d(P) = 1$, then $P = \sum_{i=1}^k P_i$, where each $P_i$ is a segment. This contradicts Lemma \ref{lem:subgraph}, which implies that at least one $P_i$ must include $k$ vertices. Therefore, we conclude $d(P) \geq 2$.

  For the sake of induction, let's assume that the result holds for all cases up to $k-1$. Now, consider that $G(P)$ includes a complete subgraph consisting of $k$ vertices. By definition, we can express $P$ as
  \[
    P = \sum_{i=1}^q \conv{P_i, Q_i}, \quad \textrm{where } d(P_i), d(Q_i) < d(P). 
  \]
  According to Lemma \ref{lem:subgraph}, there exists an index $i$ for which $G(\conv{P_i, Q_i})$ also contains a complete subgraph $K$ with $k$ vertices. Without loss of generality, we can assume that $P_i$ contains at least $\lceil \frac{k}{2}\rceil$ vertices of $K$.

  If two vertices $x, y$ of $\conv{P_i, Q_i}$ span an edge, then they also span an edge in $P_i$ if $x, y \in P_i$. Hence $G(P_i)$ contains not only at least $\lceil \frac{k}{2}\rceil$ vertices of $K$ but also the complete subgraph induced by them. By the induction hypothesis we obtain
  \[
    d(P) - 1 \geq d(P_i) \geq \Big\lceil \log_2 \Big\lceil \frac{k}{2} \Big\rceil \Big\rceil = \big\lceil \log_2 k \big\rceil - 1,
  \]
  from which it follows $d(P) \geq \lceil \log_2 k \rceil$.
\end{proof}

\section{Depth of Polytopes} \label{sec:polytopes}

We begin by computing the depth complexity of polygons, whose depth is at most 2.\\

\noindent\textbf{Polygons.} $d(P) = 1$ if $P$ is a zonotope, otherwise $d(P) = 2$.  
\begin{proof}
  If $P$ is a zonotope, then $d(P)=1$. If $P$ is a triangle, then $d(P)=2$ by \cref{prop:f}(a), since a triangle is not a zonotope.

  Now suppose $P$ is neither a zonotope nor a triangle. Then, $P$ can be decomposed as $P = \sum_{i=1}^k P_i,$ where each $P_i$ is either a zonotope or a triangle \cite[Chapter 15.1]{grunbaum2003convex}. By Proposition \ref{prop:basic_upper}(a), we obtain $d(P) = 2$. 
\end{proof}

A \textit{pyramid} is a polytope of the form $P=\conv{Q,x}$, where $Q$ is the \emph{basis} and $x\in\R^n$ such that $x\not\in\text{aff }Q$, the affine hull of $Q$.\\

\noindent\textbf{Pyramids.} $d(Q) \leq d(P) \leq d(Q) + 1$, where $Q$ is a basis of $P$.  
\begin{proof}
  From \cref{prop:basic_upper}(b) we obtain $d(P) \leq d(Q)+1$. Since $Q$ is a face of $P$, \cref{prop:faces} gives $d(Q) \leq d(P)$. 
\end{proof}

A \textit{bipyramid} is a polytope of the form $P = \conv{Q, x_1, x_2}$, where $Q$ (with $\dim Q \geq 1$) is the \textit{base}, and $x_1, x_2 \notin \text{aff }Q$ are the \textit{apices}, such that the segment $[x_1, x_2]$ intersects the relative interior of $Q$.\\

\noindent\textbf{Bipyramids.} $d(Q) \leq d(P) \leq d(Q) + 1$, where $Q$ is the basis of $P$.  
\begin{proof}
  The upper bound $d(P) \leq d(Q)+1$ follows from \cref{prop:basic_upper}(b). For the lower bound, let $H$ be a hyperplane containing aff $Q$ but not $x_1,x_2$. Projecting $P$ onto $H$ along the line through the apices gives $Q$, hence by \cref{prop:basic_upper}(c) we have $d(Q) \leq d(P)$.
\end{proof}

\noindent\textbf{Cartesian product.} $d(P_1\times P_2)=\maxb{d(P_1),d(P_2)}.$
\begin{proof}
Let
  \[
    P_1' = P_1\times\{0\}, \qquad P_2' = \{0\}\times P_2.
  \]
  Then $P_1\times P_2 = P_1' + P_2'$. Since $P_1'$ and $P_2'$ are affinely equivalent to $P_1$ and $P_2$, respectively, \cref{prop:basic_upper}(a, c) gives
  \[
    d(P_1\times P_2) \leq \maxb{d(P_1),d(P_2)}.
  \]

  For the reverse inequality, choose vertices $x_i$ of $P_i$. Then $P_1\times\{x_2\}$ and $\{x_1\}\times P_2$ are faces of $P_1\times P_2$ affinely equivalent to $P_1$ and $P_2$. By \cref{prop:faces},
  \[
    d(P_i) \leq d(P_1\times P_2),\ \  i = 1, 2.
  \]
  Hence
  \[
    \maxb{d(P_1),d(P_2)} \leq d(P_1\times P_2),
  \]
  which proves the claim.
\end{proof}

A \textit{prism} is a polytope of the form $P = \conv{Q, x + Q}$, where $Q$ is called the base of $P$ and the translation $x + Q\notin\text{aff }Q$.\\

\noindent\textbf{Prisms.} $d(P) = d(Q)$, where $Q$ is the base of $P$.  
\begin{proof} 
  A prism with base $Q$ can be written as $P=Q\times [0,1]$ up to affine equivalence. The claim follows from the cartesian product depth, since $d([0,1])=1$.
\end{proof}

The \textit{cross-polytope} is polytope of the form $P = \conv{\pm e_1, \ldots, \pm e_n}$, where $e_i$ are the standard basis vectors of $\R^n$.\\

\noindent\textbf{Cross-polytopes.} $d(P)=\lceil \log_2 n \rceil$ for $n\geq 2$.  
\begin{proof}
  The vertices $e_1,\ldots,e_n$ form a clique in the graph of $P$, so by \cref{thm:graph},
  \[
    d(P) \geq \lceil \log_2 n \rceil.
  \]

  For the upper bound, proceed by induction on $n$. The case $n=2$ is a square, hence a zonotope, so $d(P)=1$.

  Assume $n\geq 3$ and write $n=a+b$ with $a=\lfloor n/2 \rfloor$ and $b=\lceil n/2 \rceil$. Identifying $\R^n = \R^a\times\R^b$, we have
  \[
    P = \conv{P_a\times\{0\},\; \{0\}\times P_b},
  \]
  where $P_a, P_b$ are the cross-polytopes of dimension $a$ and $b$, respectively. Therefore, by \cref{prop:basic_upper}(b, c), the cartesian product depth, and the induction hypothesis,
  \begin{align*}
    d(P)
    &\leq \maxb{d(P_a\times\{0\}), d(\{0\}\times P_b)} + 1 \\
    &= \maxb{d(P_a), d(P_b)} + 1 \\
    &= \lceil \log_2 b \rceil + 1 \\
    &\leq \lceil \log_2 n \rceil.
  \end{align*}
  Hence $d(P) = \lceil \log_2 n \rceil$.
\end{proof}

A \textit{2-neighborly polytope} $P$ is a polytope where the segment between any two vertices is an edge of $P$.\\

\noindent\textbf{2-neighborly polytopes.} $d(P) = \lceil \log_2 k \rceil$, where $k$ is the number of vertices of $P$.  
\begin{proof}
  Since $G(P)$ is complete, this follows from \cref{prop:f}(a) and \cref{thm:graph}.  
\end{proof}

As simplices are 2-neighborly we obtain directly their depth complexity.\\

\noindent\textbf{Simplices.} $d(P) = \lceil\log_2 (n + 1) \rceil$.\\

Using \cref{thm:cpwl-poly}, we obtain a geometric proof of \cref{thm:arora} for ReLU networks: the Newton polytope of $\maxb{x_1, \ldots, x_n, 0}$ is an $n$-simplex, so the computed depth for simplices recovers the same depth bound as \cref{thm:arora}.

The \textit{join} of polytopes is a polytope of the form $P_1 * \cdots * P_k = \conv{P_1, \ldots, P_k}$, where the $P_i$ lie in pairwise skew affine subspaces.\\

\noindent\textbf{Join of $(2^{m_i}-1)$-simplices.} $d(P) = \Big\lceil \log_2\Big(\sum_{i=1}^k 2^{m_i}\Big)\Big\rceil$, for $P = P_1 * \cdots * P_k$, where each $P_i$ is a $(2^{m_i}-1)$-simplex.
\begin{proof}
  For each $i$, the simplex $P_i$ has depth $d(P_i)=m_i$. Since $P = \conv{P_1, \ldots, P_k}$, \cref{prop:optimal_hull} gives
  \[
    d(P) \leq \Big\lceil \log_2\Big(\sum_{i=1}^k 2^{m_i}\Big)\Big\rceil.
  \]

  On the other hand, the graph of $P$ is complete on $\sum_{i=1}^k 2^{m_i}$ vertices, since every pair of vertices from different $P_i$ is adjacent. Therefore, \cref{thm:graph} yields
  \[
    d(P) \geq \Big\lceil \log_2\Big(\sum_{i=1}^k 2^{m_i}\Big)\Big\rceil,
  \]
  proving the claim.
\end{proof}

The depth computation for join of simplices shows that for every vector $(m_1, \ldots, m_k) \in \mathbb{N}^k$, there are polytopes attaining exactly the upper bound of \cref{prop:optimal_hull}.

A \textit{cyclic polytope} $C_n(k), k > n,$ is a polytope of the form
\[
  C_n(k) = \conv{\gamma(t_1), \ldots, \gamma(t_k)},
\]
where $t_1 < \ldots < t_k$ are real numbers and $\gamma(t) = (t, t^2, \ldots, t^n)$. For $n\geq 4$, cyclic polytopes are 2-neighborly, therefore we obtain directly their depth complexity.\\

\noindent\textbf{Cyclic polytopes.}  $d(C_n(k)) = \lceil\log_2 k \rceil$ for $n\geq 4$.\\

Since $d(P) \leq d_0(P)$, where $d_0(P)$ is the minimum ICNN depth required to represent $\mathcal{F} P$, the depth of cyclic polytopes for $n\geq 4$ shows that ICNNs do not satisfy a result like \cref{thm:arora}. Unlike general ReLU networks, no fixed depth suffices to represent $\mathcal{F} C_n(k)$ as $k$ grows.

For $n=2$, polygons were shown to have bounded depth $d(P)\leq 2$, while the case $n=3$ remains open. We show that bipyramids with a triangular base provide examples of polyhedra of depth 3. This demonstrates that depth complexity in dimension $n=3$ behaves differently from the case $n=2$.

Before computing the depth complexity of triangular bipyramids, we first introduce a simple tool for indecomposable polytopes.

Two polytopes, $P$ and $Q$, are said to be \textit{positively homothetic}, if $P = \lambda Q + w$ for some $\lambda > 0$ and $w\in\R^n$. A polytope $P$ is said to be \textit{indecomposable} if any decomposition $P = \sum_{i=1}^k P_i$ is only possible when $P_i$ is positively homothetic to $P$ for all $i=1, \ldots, k$. 

\begin{lemma} \label{lem:indecomposable}
  If $P$ is an indecomposable polytope, then there exist polytopes $P_1, P_2$ such that $P = \conv{P_1, P_2}$ and $d(P) = \maxb{d(P_1), d(P_2)} + 1$.
\end{lemma}
\begin{proof}
  By definition, if $d(P) = m$, then $P$ can be written as
  \[
    P = \sum_{i = 1}^q \conv{P_{i1}, P_{i2}}, \quad d(P_{ij}) < m \quad \forall i,j,
  \]
  where $m$ is the smallest integer for which such a representation exists. Hence there must exist an index $i$ such that $\maxb{d(P_{i1}), d(P_{i2})} + 1 = d(P)$.

  Since $P$ is indecomposable, there are $\lambda_i > 0$ and $w_i \in \mathbb{R}^n$ where $P$ satisfies
  \[
    P = \lambda_i \conv{P_{i1}, P_{i2}} + w_i = \conv{\lambda_i P_{i1} + w_i, \lambda_i P_{i2} + w_i}.
  \]
  Each $\lambda_i P_{ij} + w_i$ is an affine image of $P_{ij}$ under an invertible map. By \cref{prop:basic_upper}(c), affine equivalence preserves depth, so 
  \[
    d(P) = \maxb{d(P_{i1}), d(P_{i2})} + 1 = \maxb{d(\lambda_i P_{i1} + w_i), d(\lambda_i P_{i2} + w_i)} + 1. \qedhere
  \]
\end{proof}

\noindent\textbf{Triangular bipyramids in $\R^3$.} $d(P) = 3$.  
\begin{proof}
  A bipyramid with a 2-depth basis satisfies $2 \leq d(P) \leq 3$. Assume, for contradiction, that $d(P)=2$. As $P$ is indecomposable \cite[Chapter 15.1]{grunbaum2003convex}, Lemma \ref{lem:indecomposable} implies that there exist $P_1, P_2$ such that $P=\conv{P_1,P_2}$ and $\max\{d(P_1),d(P_2)\} + 1=2$. Without loss of generality, suppose $P_1$ is a zonotope containing at least three vertices of $P$.

  There are three cases: $P_1$ contains two base vertices and an apex; or the three base vertices; or one base vertex and the two apices. In each case $P_1$ is not centrally symmetric, contradicting that it is a zonotope. Hence $d(P)\neq 2$ and so $d(P)=3$. 
\end{proof}

We conclude with a complementary construction: whereas for cyclic polytopes the depth grows with the number of vertices, here we obtain polytopes with arbitrarily many vertices at fixed depth.

\begin{theorem} \label{thm:fixed}
  For any $n \geq 5$ and $m\in\N$, there exists a family of $n$-polytopes with arbitrarily many vertices and depth $m$.
\end{theorem}
\begin{proof}
  Let $g_i = [\mathbf{0}, b_i], i = 1, \ldots, p$, represent line segments from the origin with $b_1, \ldots, b_p$ denoting $p \geq n$ points in general position. The zonotope $Z_p = \sum_{i=1}^p g_i$ has depth 1 and $v_p = 2\sum_{i=0}^{n-1}\binom{p-1}{i}$ vertices \cite{zaslavsky1975facing}, which grows with $p$.

  For any $m \geq 2$, let $H$ be a supporting hyperplane of a vertex $x$ of $Z_p$, and take a polytope $P \subset H$ with $\dim P < n$ and $d(P)=m$ (e.g. a simplex or cyclic polytope). 

  Then $P+Z_p$ satisfies $d(P+Z_p) \leq m$ by \cref{prop:basic_upper}(a). On the other hand, the face of $P+Z_p$ in the direction of $x$ is a translate of $P$, hence by \cref{prop:faces}, $m = d(P) \leq d(P+Z_p)$. Thus $P + Z_p$ forms a $m$-depth family whose number of vertices increases with $p$.
\end{proof}

Similar constructions exist for $n \leq 4$ and $m \leq 3$, but a full generalization of Theorem \ref{thm:fixed} for $n=3,4$ and arbitrary $m$ would require a better understanding of depth complexity in dimension $3$.

\bibliographystyle{plain}
\bibliography{bib}

\end{document}